OpenStudy (shesolitt):

PLEASE HELP WILL MEDAL AND FAN! Given the equation 7(x + 2) = 2x − 1, solve for the variable. Explain each step and justify your process. B. Jeff solved a similar equation below. Is Jeff's solution correct? Explain why or why not. 3x − 2 = 5(x + 2) 3x − 2 = 5x + 10 8x = 8 x = 1

OpenStudy (mayankdevnani):

A) \[\large \bf 7(x+2)=2x-1\] apply distributive property, \[\large \bf 7x+14=2x-1\] subtract 2x from both the sides, \[\large \bf5x+14=-1\] subtract 14 from both the sides, \[\large \bf 5x=-15\] divide 5 from both the sides, \[\large \bf x=-3\]

OpenStudy (mayankdevnani):

similarly , solve B) part

OpenStudy (acespeedfighter):

how did you get 2 answers?

OpenStudy (acespeedfighter):

well questions i mean

OpenStudy (shesolitt):

what do you mean? lol @AceSpeedFighter I didn't turn this assignment in yet

OpenStudy (pythagoras123):

B) Jeff's solution is incorrect. The equation 3x-2 = 5x+10 Simplifies to -2x = 12 x = -6.

OpenStudy (shesolitt):

so many people are saying different things

OpenStudy (elusive):

Should have done 3x MINUS 5x instead of plus. and 10 PLUS 2 instead of minus.

OpenStudy (shesolitt):

thank you sm. I'm gonna close this question can you help me with a few more @elusive

OpenStudy (elusive):

No problem